3. Let X be a random variable and denote by Mx(t) its MGF. Prove that, for any t > 0, we have P[X >Mx(t)e
Answer #1

Markov inequality states that, if Y is a non negative random variable,

for a>0,

P(Y>a)\leq \frac{E(Y)}{a}

Note that, e^{tX} is non negative.

So,

P(X>1)=P(tX>t)=P(e^{tX}>e^t)\leq \frac{E(e^{tX})}{e^t} (by markov inequality)

where,  Y=e^{tX}, a=t, M_X(t)=E(e^{tX})

So,

Mx t) P(X > 1) P(X >1) 3 Mx(t)e